OpenStudy (anonymous):

Xavier is riding on a Ferris wheel at the local fair. His height can be modeled by the equation H(t) = 20 cospi over 15 + 30, where H represents the height of the person above the ground in feet at t seconds.

OpenStudy (anonymous):

Part 2: How long does the Ferris wheel take to make one complete revolution? Part 3: Assuming Xavier begins the ride at the top, how far from the ground is the edge of the Ferris wheel, when Xavier's height above the ground reaches a minimum?

OpenStudy (anonymous):

part 1 I believe is 28

OpenStudy (anonymous):

Given equation follows the form: y=Acos(Bx-C)+D, A=amplitude, period=2π/B, C/B=phase shift, D=vertical shift

OpenStudy (anonymous):

thanks :) part 3?

OpenStudy (anonymous):

that was for part 1

OpenStudy (anonymous):

I know. But like 28 seconds?

OpenStudy (anonymous):

what about the next part.

OpenStudy (anonymous):

part 1 28 ft

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

frankly speaking don't know the answer to part 2 but just give me some time i will try my best

OpenStudy (anonymous):

yup don't worry! I appreciate your help

OpenStudy (anonymous):

ok can u tell me the period for part 2 by looking at the information i have provided u (Given equation follows the form: y=Acos(Bx-C)+D, A=amplitude, period=2π/B, C/B=phase shift, D=vertical shif)

OpenStudy (anonymous):

ummm no i dont know it :(

OpenStudy (anonymous):

i suppose the given equation is H(t) = 20 cospi over 15 + 30

OpenStudy (anonymous):

okay! true

OpenStudy (anonymous):

H(t)=y

OpenStudy (anonymous):

How would i solve that

OpenStudy (anonymous):

u dont have to solve that

OpenStudy (anonymous):

u have to substitute y in the place of H(t) in H(t) = 20 cospi over 15 + 30

OpenStudy (anonymous):

But how will I get the answer to the last part...

OpenStudy (anonymous):

Assuming Xavier begins the ride at the top, how far from the ground is the edge of the Ferris wheel, when Xavier's height above the ground reaches a minimum?

OpenStudy (anonymous):

y=Acos(Bx-C)+D

OpenStudy (anonymous):

I have to answer that!

OpenStudy (anonymous):

Can you help me?

OpenStudy (anonymous):

u have to find the period for that

OpenStudy (anonymous):

period=2π/B

OpenStudy (anonymous):

the equation is y= 20 cospi over 15 + 30

OpenStudy (anonymous):

yesss but how would i solve it :( i need to find the minimum

OpenStudy (anonymous):

Given equation follows the form: y=Acos(Bx-C)+D

OpenStudy (anonymous):

i am really sorry but i can't the information i provided with in this question was by taking the help of internet as i am not familiar with this chapter itself i tried my best sry

OpenStudy (anonymous):

whatever its fine

OpenStudy (anonymous):

but yes i can provide u with the answer of somewhat a similar quaetion

OpenStudy (anonymous):

Travis is riding the Ferris wheel at the amusement park. His height can be modeled by the equation H(t) = 22 cos (pi over 13)t + 28, where H represents the height of the person above the ground in feet at t seconds. *** Given equation follows the form: y=Acos(Bx-C)+D, A=amplitude, period=2π/B, C/B=phase shift, D=vertical shift .. For given equation: H(t) = 22 cos (pi over 13)t + 28 amplitude=22 B=π/13 period=2π/B=2π/(π/13)=26 sec Phase shift=0 vertical shift=28 ... Part 1: How far above the ground is Travis before the ride begins? H(0) = 22 cos (pi over 13)*0 + 28=22+58=50 ft H(0) = 22 cos (0) + 28=22+58=50 ft .. Part 2: How long does the Ferris wheel take to make one complete revolution? period=2π/B=2π/(π/13)=26 sec .. Part 3: Assuming Travis begins the ride at the top, how far from the ground is the edge of the Ferris wheel, when Travis' height above the ground reaches a minimum? Cos function reaches a minimum at 1/2 period=13 sec H(13)=22cos(π/13*13)+vertical shift of 28 up H(13)=22cos(π)+vertical shift of 28 H(13)=22*(-1)+ 28=-22+28=6 ft

OpenStudy (anonymous):

the only difference is that some of its values are different

OpenStudy (anonymous):

was that helpful?